OpenStudy (anonymous):

Find the solutions to the equation (check all that apply) 20x2-26x+8=0. a.x=1/5 b.x=1/2 c.x=4/5 d.x=2/3 e.x=1/3 f.x=3/5

OpenStudy (anonymous):

Help please!

OpenStudy (radar):

x=4/5 and x=1/2

OpenStudy (radar):

Use quadratic formula.

OpenStudy (anonymous):

20x^2-26x+8=0 20x^2-10x-16x+8=0 [20*8=160; 160=10*16; -10-16=-26 ] 10x(2x-1)-8(2x-1)=0 (10x-8)(2x-1)=0 x=1/2 and x=4/5 [ANS]